Delectable Parsnip, Inc.’s, net income for the most recent year was $8,795. The tax rate was 21 percent. The firm paid $4,653 in total interest expense and deducted $5,124 in depreciation expense. What was the company’s cash coverage ratio for the year?
Answer #1

(EBIT-interest expense)(1-tax rate)=Net income

(EBIT-4653)(1-0.21)=8,795

EBIT-4653=8,795/0.79

EBIT=(8,795/0.79)+4653

=$15785.9114

Cash coverage ratio=(EBIT+depreciation)/interest expense

=(15785.9114+5124)/4,653

=4.49(Approx).